40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10215     Entry time: Wed Jul 16 07:51:52 2014
Author: Akhil 
Type: Update 
Category: General 
Subject: Weekly Update 

Work Done:

  • Solved all the timing issues pertaining to the R Pi and the FC.
  • Took all the measurements for complete characterization of the frequency counter(Phase Plots to follow shortly).
  • Finished  installation of the FC on the martian and created a channel  for the FC frequencies(will be tested in this week).

Plans for this Week:

  • Testing of the EPICS soft IOC created for the FC as a channel access server and hence completing the installation of the FC.
  • Placing the FC inside the lab( plan discussed in this elog: http://nodus.ligo.caltech.edu:8080/40m/10163) with proper supervision.
  • Characterization of the temperature actuator.

Inside the 40m Lab:        

I will need to be inside the lab to place the FC . This will be done in the morning session (on thursday) with supervision of Manasa and Steve(if required).

 

 

 

 

 

 

 

 

ELOG V3.1.3-